I would like to know whether the answer I have chosen is correct or incorrect.

My answer: A

Question:

3. Given the following information for the Network Company: Date December 31, 2004 December 31,2005 December 31, 2006 Market $ 800 940 1,060 Cost 800 1,000 1,100 If the allowance method of recording lower of cost or market is in use, which December 31, 2006 entry is not correct? 40 20 20 20 Loss Due toMarket Valuation 40 20 20 20 Allowance to Reduce Inventory to Market b. Allowance to Reduce Inventory to Market c. Allowance to Reduce Inventory to Market d. Allowance to Reduce Inventory to Market Loss Recovery Due to Market Valuation Gain on Revaluation of Inventory to Market Cost of Goods Sold Note: b. is also correct.
Answer #1

correct option is "D"

The question says which of the following entry is not correct.

Why other option is correct

Option 1)since the market value is lower than cost ,loss due to market valuation will be debited and allowance to reduce inventory to market will be credited by [1100-1060] 40

Option 2)At december 2004 ,since cost is equal to market value ,no loss is recorded.

At december 2005 ,since market value is lower than cost by 1000-940=60 (entry made :,loss due to market valuation will be debited and allowance to reduce inventory to market will be credited by 60]

At december 2006 ,since market value is lower than cost by 1100-1060=40 (however there is already a balance in allowance account amounting to 60 made in 2005 ,so allowance for 20 will be reversed (60-40) since excess]

option C)At december 2006 ,since market value is lower than cost by 1100-1060=40 (however there is already a balance in allowance account amounting to 60 made in 2005 ,so gain due to revaluation of inventory will be made by 20 ]

This 3 enteries are correct only entry incorrect is "D"